Scope = 10chairs @ $10.00-$100.00 Timeframe 10 days At the End of the 5th day 4 chairs were Completed at a cost of $11.00 each PV: AC EV CPI- sv=
Answer #1

Project ABC: Assume that each chair cost $10 and take one day to complete

Planned Value (PV) is the value which was supposed to earn at that point of time

Planned value (PV) calculation at the end of 5th day

PV = (% planned completion) * Budgeted Cost

= 50% * $100 = $50

Earned Value (EV) is valued earned to the proportion of the completion of the task at point of time

EV = (% completion) * Budgeted Cost

= 40% * $100 = $40

The actual cost (AC) of project = completed chairs * Actual Cost

= 4 *$11 = $44

Cost variance (CV) = EV – AC

= $40 - $44 = -$4

Cost performance index               (CPI) = EV/AC = $40 /$44 = 0.91

Schedule Variance (SV) = Earned Value (EV) - Planned Value (PV)

= $40 - $50 = -$10

Schedule performance index (SPI) = EV/PV = $40/$50 = 0.8

Critical Ratio (CR) is a product of the cost performance index (CPI) and schedule performance index (SPI)

CR = CPI * SPI = 0.91 * 0.8 = 0.73